Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

florist (probability)

Expert replies
by rahul.s » Mon Feb 15, 2010 11:09 pm
a florist has 2 roses, 3 buttercups, and 4 orchids. she puts 2 flowers together at random in a bouquet. however, the customer calls and says that she does not want 2 of the same flower. what is the probability that the florist does not have to change the bouquet?

sorry, but there aren't any answer choices.

OA: [spoiler]13/18[/spoiler]
Source: mgmat
Join the discussion
Source: — Problem Solving |

by ajith » Mon Feb 15, 2010 11:41 pm
rahul.s wrote:a florist has 2 roses, 3 buttercups, and 4 orchids. she puts 2 flowers together at random in a bouquet. however, the customer calls and says that she does not want 2 of the same flower. what is the probability that the florist does not have to change the bouquet?

sorry, but there aren't any answer choices.

OA: [spoiler]13/18[/spoiler]
Source: mgmat
No of ways in which one can select two flowers from 9 (3+4+2) = 9C2 = 36
No of ways in which one can select two flowers of the same kind = 2C2 + 3C2+ 4C2 = 1+3+6 =10
Probability that the two flowers that selected that of same kind = 10/36 = 5/18
Probability that the florist does not have to change the bouquet = 1 -Probability that the two flowers that selected that of same kind = 1-5/18 = 13/18
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by thephoenix » Tue Feb 16, 2010 12:19 am
tot no. of wyas for selecting 2 flower out of 9 flower=9C2=36

lets find out the cases where the two flowers are same
P of two flowers to be roses=2C2/9C2=1/36
P of two flowers to be buttecups=3C2/9C2=3/36
P of two flowers to be buttecups=4C2/9C2=6/36

tot P=add all=10/36=5/18

reqd P=1-tot P for selceting two of same kind
=13/18
Join the discussion

by kvamsy » Tue Feb 16, 2010 5:26 am
rahul.s wrote:a florist has 2 roses, 3 buttercups, and 4 orchids. she puts 2 flowers together at random in a bouquet. however, the customer calls and says that she does not want 2 of the same flower. what is the probability that the florist does not have to change the bouquet?

sorry, but there aren't any answer choices.

OA: [spoiler]13/18[/spoiler]
Source: mgmat
One more way to solve.

Total number of ways of selecting 2 flowers = 9C2 = 36
2 flowers should select at random = (2C1*3C1)+(2C1*4C1)+(3C1*4C1)

probability = 26/36 -----> 13/26
Join the discussion

by rahul.s » Tue Feb 16, 2010 5:34 am
kvamsy wrote:One more way to solve.

Total number of ways of selecting 2 flowers = 9C2 = 36
2 flowers should select at random = (2C1*3C1)+(2C1*4C1)+(3C1*4C1)

probability = 26/36 -----> 13/26
i think you meant 13/18 :)
Join the discussion

by kvamsy » Tue Feb 16, 2010 5:51 am
rahul.s wrote:
kvamsy wrote:One more way to solve.

Total number of ways of selecting 2 flowers = 9C2 = 36
2 flowers should select at random = (2C1*3C1)+(2C1*4C1)+(3C1*4C1)

probability = 26/36 -----> 13/26
i think you meant 13/18 :)
you are right
Join the discussion